# Mixture of Directed Graphical Models for Discrete Spatial Random Fields

### v3

@@ Appendix C Conditional Independence Relationships in Graphs and Probability Distributions
As stated in the introduction, an MRF is a probability distribution whose conditional distributions have the same dependence structure as the NUG. In the terms of conditional independence relationship sets, we can alternatively define an MRF, $p(\mathbf{z})$, specified for a NUG, $\mathbf{N}$, as a probability distribution such that $I(p)=I(\mathbf{N})$ is true. We previously used the local Markov property, that is a vertex, $i$, is independent of all other vertices given its neighbors, to obtain the full conditionals of $z_{i}$ given the set of neighbors $z_{\partial(i)}$ (Equation 6). Additional conditional independence relationships are also asserted by the NUG. The most general rule to determine whether $\mathbf{A}$ and $\mathbf{B}$ are independent given $\mathbf{C}$ is if all paths from $\mathbf{A}$ to $\mathbf{B}$ go through the set $\mathbf{C}$. This is known as the global Markov property. From this definition of independence – defined through the graphical representation of dependence between variables – follows the local Markov property and an additional property known as pairwise Markov property. The pairwise Markov property states that any two nodes $i$ and $j$ are independent given all other vertices if $i$ and $j$ are not neighbors. In Figure 5, $A\perp E|\{C,B,D,F\}$, as there is not an edge between $A$ and $E$ and all paths connecting $A$ and $E$ go through the conditioning set. Again in Figure 5, we can verify that $B\perp\{C,F\}|\{A,D,E\}$ as there is no path from $B$ to $C$ or $B$ to $F$ that does not go through $A,D$ or $E$, the neighbors of $B$. We can prove the pairwise and local properties from the global property and, incidentally, we can use any one of the properties to prove the other two when the probability distribution over the state space is strictly positive, highlighting the intuitive power of the graphical representation (Koller and Friedman, 2009, pg 119).
@@
While the Markov property for DAGs is intuitive and leads to the convenient factorization of the corresponding DGM, more general statements about the set of conditional independence relationships in a DAG, $I(\mathbf{D})$, require the concept of directed separation (d-separation). Recall that a path is a sequence of undirected and connected edges. A path in a DAG is then a sequence of connected edges where the direction is ignored. There are two rules that govern d-separation for sets of variables $\mathbf{A}$ and $\mathbf{B}$ given $\mathbf{C}$, that is conditional independence statements of the form $\mathbf{A}\perp\mathbf{B}|\mathbf{C}$. We say that $\mathbf{A}$ and $\mathbf{B}$ are separated by $\mathbf{C}$ if all the paths from $\mathbf{A}$ to $\mathbf{B}$ together imply independence when conditioned on $\mathbf{C}$. The rules of d-separation depend on the notion of a collider. A collider vertex at which two directed edges meet head to head. In Figure 6 the vertices $B$ and $D$ are both colliders. The set of vertices that have directed edges oriented towards the collider is called a v-structure. $A\rightarrow B\leftarrow E$, is a v-structure in Figure 6. For a path that contains a collider, the path separates two sets of variables $\mathbf{A}$ and $\mathbf{B}$, if the collider or any the descendants of the collider are not in the set $\mathbf{C}$. Second, $\mathbf{A}$ and $\mathbf{B}$ are separated by $\mathbf{C}$ if every path that does not contain a collider has a vertex that is in $\mathbf{C}$. Using these two rules we can see that $A\perp E|C$, as the path $A-C-E$ does not contain a collider and we condition on $C$, whereas the path $A-B-E$ does contain a collider and $B$ is not conditioned on. We can also assert that $A\not\perp E|C,B$, since the path with the collider $B$ is now in the conditioning set. Additionally, $A\not\perp E|C,D$ and $A\not\perp E|C,F$, since $D$ and $F$ are descendants of the collider $B$.
C.1 Comparison of Conditional Independence Relationships
@@
Now we further explore the differences between $I(\mathbf{D})$ and $I(\mathbf{N})$ and consider the common case where $I(\mathbf{D})\neq I(\mathbf{N})$, nor is $I(\mathbf{D})\not\subset I(\mathbf{N})$, nor $I(\mathbf{N})\not\subset I(\mathbf{D})$. That is, a NUG and DAG cannot be used to represent the same probability distribution, neither perfectly nor as an I-map. Consider the example of the NUG in Figure 7(a) which asserts that $B\perp D|A,C$ but $B\not\perp D$. We can create a DAG from Figure 7(a) by assigning an acyclic orientation to the vertices. In doing so, we will invariably change the set of conditional independence relationships that are asserted by the original undirected graph. Figures 7(b) and 7(c) are two failed attempts to construct a DAG with the same independence relationships as the undirected graph in Figure 7(a). In Figure 7(b), $B\perp D$ but $B\not\perp D|A,C$, the complement of the two independence relationships we stated about Figure 7(a). Figure 7(c) correctly asserts that $B\not\perp D$, however, shows that $B\not\perp D|A,C$ which is incorrect for Figure 7(a). In fact, there is no DAG that can encode the same conditional independence relationships as the given NUG. Any DAG that we create from the NUG by assigning a direction to each edge in the graph will invariably result in a v-structure, which introduces dependence between previously unconnected vertices when the collider is in the conditioning set. The introduction of v-structures in the the most obvious violation of the pairwise Markov property when creating a DAG from a NUG.
@@
There are special cases, however, where $I(\mathbf{D})=I(\mathbf{N})$, when the skeleton of $\mathbf{D}$ is also equivalent to $\mathbf{N}$. The skeleton of a DAG $\mathbf{D}$, is the undirected graph created as $\mathbf{U}(\mathbf{D})=\{\mathbf{V},\{\{i,j\}:(i,j)\in\vec{\mathbf{E}}\}\}$. When two graphs, $\mathbf{G}$ and $\mathbf{G}^{\prime}$, represent the same conditional independence relationships, $I(\mathbf{G})=I(\mathbf{G})$, they are said to be independence equivalent (I-equivalent). Chordal graphs are a special case of graph for which a DAG and an undirected graph can represent the equivalent sets of conditional independence relations. A chordal graph is a graph in which every undirected cycle of four or more vertices has edges which are not part of the cycle but connect vertices within the cycle. The graph in Figure 7(a) is nonchordal, as there are no edges connecting the four vertices of the cycle that are not part of the cycle. Figure 7(a) could be made into chordal graph by adding an edge from either $A$ to $C$, $B$ to $D$, or both. Any DAG created from a chordal NUG, $\mathbf{G}$, by assigning a direction to each edge in $\mathbf{G}$ and does not introduce v-structures for vertices whose parents are not already connected, will represent an equivalent set of independence relationships as asserted by $\mathbf{G}$. By definition of chordality, in chordal graphs it is always possible to find an acyclic orientation in which the parents of a collider, as defined by the acyclic orientation, are already connected. The NUG, $\mathbf{N}$, is said to be an equivalence class of all DAGs generated by assigning an acyclic orientations to $\mathbf{N}$ which do not introduce v-structures for unconnected parents (He et al., 2015). While practically zero areal data analysis problems admit a chordal graph to define the neighborhood structure between areal units, as referenced in Section 3.1, we utilized this relationship to our advantage when subgraphs of the NUG are chordal in order to create DAGs which are Markov blanket equivalent with the NUG. We also used this relationship to our computational advantage for tree graphs, another special instance of a chordal graph. Since a tree is a chordal graph, the undirected graph is an equivalence class for all DAGs with the same skeleton in which all edges are oriented away from a single vertex, called the root. Again, it can be shown that specifying a probability distribution for a tree graph can be written equivalently as an MRF or as a DGM (Meila and Jordan, 2001).
@@
In this Section, we have only compared the conditional independence assumptions of a single compatible DAG and NUG. The set of conditional independence relationships implied by the collection of DAGs used in the mixture model to specify the DGM components differs from that of a single DAG.
Appendix D Simulation Study
D.1 Phase Transition
@@
D.2 Complete Data Setting
In the complete data scenario, we set $m_{i}=2$ for all $i=1,\dots,n$ and perform the simulation across all combinations of $\beta=\{0.1,0.15,0.2,0.25,0.3\}$ and $\eta=\{0.05,0.2\}$. The results are shown in Figure 8.
@@ D.2 Complete Data Setting
Consistent with the missing data regime, we observe similar performance of all five models in the low uncertainty setting (i.e. $\eta=0.05$). Likewise in the high uncertainty setting, the aMRF has relative worse performance in estimating the true clustering of the latent field as the spatial dependence parameter increases. The AO and rooted MDGM priors show a relative increase in performance as $\beta$ increases and the MDGM-ST consistently matches the performance of the MRF in estimating the true dependence of the latent field.
Appendix E Cross Validation Study
For each iteration of the cross validation study we randomly select 60 areal units with ratings to be withheld as the test set. Using the remaining ratings as our training set, we fit the model with both a MDGM-ST and aMRF prior on the latent variable. As the withheld data is the observed ratings, we compare using the mean absolute difference of the posterior mean predicted probability of a block group having a rating of one to the average observed ratings at the block group. We take the average of the absolute errors for the block groups in the test set as our summary statistic of model performance for a single iteration of the study. We performed one hundred iterations of the cross validation study. We found that the mean absolute error of the posterior predicted probability of a rating of one to the average observed ratings for the model with an MDGM-ST prior is .2188, and the mean absolute error for the aMRF prior is .2204. The maximum absolute difference between the mean absolute errors of the two models across iterations was .007, indicating a practically equivalent performance of the two models.

### v4

@@ Appendix G Conditional Independence Relationships in Graphs and Probability Distributions
As stated in the introduction, an MRF is a probability distribution whose conditional distributions have the same dependence structure as the NUG. In terms of conditional independence relationship sets, we can alternatively define an MRF, $p(\mathbf{z})$, specified for a NUG, $\mathbf{N}$, as a probability distribution such that $I(p)=I(\mathbf{N})$ is true. We previously used the local Markov property, that is a vertex, $i$, is independent of all other vertices given its neighbors, to obtain the full conditionals of $z_{i}$ given the set of neighbors $\mathbf{z}_{\partial(i)}$ (Equation 8). Additional conditional independence relationships are also asserted by the NUG. The most general rule to determine whether $A$ and $B$ are independent given $C$ is if all paths from $A$ to $B$ go through the set $C$. This is known as the global Markov property. From this definition of independence – defined through the graphical representation of dependence between variables – follows the local Markov property and an additional property known as pairwise Markov property. The pairwise Markov property states that any two nodes $i$ and $j$ are independent given all other vertices if $i$ and $j$ are not neighbors. In Figure 5, $a\perp e|\{c,b,d,f\}$, as there is not an edge between $a$ and $e$ and all paths connecting $a$ and $e$ go through the conditioning set. Again in Figure 5, we can verify that $b\perp\{c,f\}|\{a,d,e\}$ as there is no path from $b$ to $c$ or $b$ to $f$ that does not go through $a,d$ or $e$, the neighbors of $b$. We can prove the pairwise and local properties from the global property and, incidentally, we can use any one of the properties to prove the other two when the probability distribution over the state space is strictly positive, highlighting the intuitive power of the graphical representation (Koller and Friedman, 2009, pg 119).
@@
While the Markov property for DAGs is intuitive and leads to the convenient factorization of the corresponding DGM, more general statements about the set of conditional independence relationships in a DAG, $I(\mathbf{D})$, require the concept of directed separation (d-separation). Recall that a path is a sequence of undirected and connected edges. A path in a DAG is then a sequence of connected edges where the direction is ignored. There are two rules that govern d-separation for sets of variables $A$ and $B$ given $C$, that is conditional independence statements of the form $A\perp B|C$. We say that $A$ and $B$ are separated by $C$ if all the paths from $A$ to $B$ together imply independence when conditioned on $C$. The rules of d-separation depend on the notion of a collider. A collider vertex at which two directed edges meet head to head. In Figure 6 the vertices $b$ and $d$ are both colliders. The set of vertices that have directed edges oriented towards the collider is called a v-structure. $a\rightarrow b\leftarrow e$, is a v-structure in Figure 6. For a path that contains a collider, the path separates two sets of variables $A$ and $B$, if the collider or any of the descendants of the collider are not in the set $C$. Second, $A$ and $B$ are separated by $C$ if every path that does not contain a collider has a vertex that is in $C$. Using these two rules we can see that $a\perp e|c$, as the path $a-c-e$ does not contain a collider and we condition on $c$, whereas the path $a-b-e$ does contain a collider and $b$ is not conditioned on. We can also assert that $a\not\perp e|c,b$, since the path with the collider $b$ is now in the conditioning set. Additionally, $a\not\perp e|c,d$ and $a\not\perp e|c,f$, since $d$ and $f$ are descendants of the collider $b$.
G.1 Comparison of Conditional Independence Relationships
@@
The following discussion is modified from the example used in Chapter 19 of Murphy (2012) as we further explore the differences between $I(\mathbf{D})$ and $I(\mathbf{N})$ and consider the common case where $I(\mathbf{D})\neq I(\mathbf{N})$, nor is $I(\mathbf{D})\not\subset I(\mathbf{N})$, nor $I(\mathbf{N})\not\subset I(\mathbf{D})$. That is, a NUG and DAG cannot be used to represent the same probability distribution, neither perfectly nor as an I-map. Consider the example of the NUG in Figure 7(a) which asserts that $b\perp d|a,c$ but $b\not\perp d$. We can create a DAG from Figure 7(a) by assigning an acyclic orientation to the vertices. In doing so, we will invariably change the set of conditional independence relationships that are asserted by the original undirected graph. Figures 7(b) and 7(c) are two failed attempts to construct a DAG with the same independence relationships as the undirected graph in Figure 7(a). In Figure 7(b), $b\perp d$ but $b\not\perp d|a,c$, the complement of the two independence relationships we stated about Figure 7(a). Figure 7(c) correctly asserts that $b\not\perp d$, however, shows that $b\not\perp d|a,c$ which is incorrect for Figure 7(a). In fact, there is no DAG that can encode the same conditional independence relationships as the given NUG. Any DAG that we create from the NUG by assigning a direction to each edge in the graph will invariably result in a v-structure, which introduces dependence between previously unconnected vertices when the collider is in the conditioning set. The introduction of v-structures is the most obvious violation of the pairwise Markov property when creating a DAG from a NUG.
@@
There are special cases, however, where $I(\mathbf{D})=I(\mathbf{N})$, when the skeleton of $\mathbf{D}$ is also equivalent to $\mathbf{N}$. The skeleton of a DAG $\mathbf{D}$, is the undirected graph created as $\mathbf{U}(\mathbf{D})=\{\mathbf{V},\{\{i,j\}:(i,j)\in\vec{\mathbf{E}}\}\}$. When two graphs, $\mathbf{G}$ and $\mathbf{G}^{\prime}$, represent the same conditional independence relationships, $I(\mathbf{G})=I(\mathbf{G})$, they are said to be independence equivalent (I-equivalent). Chordal graphs are a special case of graph for which a DAG and an undirected graph can represent the equivalent sets of conditional independence relations. A chordal graph is a graph in which every undirected cycle of four or more vertices has edges which are not part of the cycle but connect vertices within the cycle. The graph in Figure 7(a) is nonchordal, as there are no edges connecting the four vertices of the cycle that are not part of the cycle. Figure 7(a) could be made into chordal graph by adding an edge from either $a$ to $c$, $b$ to $d$, or both. Any DAG created from a chordal NUG, $\mathbf{G}$, by assigning a direction to each edge in $\mathbf{G}$ and does not introduce v-structures for vertices whose parents are not already connected, will represent an equivalent set of independence relationships as asserted by $\mathbf{G}$. By definition of chordality, in chordal graphs it is always possible to find an acyclic orientation in which the parents of a collider, as defined by the acyclic orientation, are already connected. The NUG, $\mathbf{N}$, is said to be an equivalence class of all DAGs generated by assigning an acyclic orientations to $\mathbf{N}$ which do not introduce v-structures for unconnected parents (He et al., 2015). While practically zero areal data analysis problems admit a chordal graph to define the neighborhood structure between areal units, as referenced in Section 3.1, we utilized this relationship to our advantage when subgraphs of the NUG are chordal in order to create DAGs which are strongly compatible with the NUG. We also used this relationship to our computational advantage for tree graphs, another special instance of a chordal graph. Since a tree is a chordal graph, the undirected graph is an equivalence class for all DAGs with the same skeleton in which all edges are oriented away from a single vertex, called the root. Again, it can be shown that specifying a probability distribution for a tree graph can be written equivalently as an MRF or as a DGM (Meila and Jordan, 2001).
@@
In this section, we have only compared the conditional independence assumptions of a single compatible DAG and NUG. The set of conditional independence relationships implied by the collection of DAGs used in the mixture model to specify the DGM components differs from that of a single DAG, which we explore below.
G.2 Conditional Independence in a Collection of DAGs
In the MDGM, the probability distribution is defined with respect to a collection of DAGs, rather than a single DAG. Importantly, if the conditional independence relationship, $A\perp B|C$ holds for every $\mathbf{D}\in\mathcal{D}$, where $\mathcal{D}$ is some set of DAGs, this does not imply that $\mathbf{z}_{A}\perp\mathbf{z}_{B}|\mathbf{z}_{C}$ holds in the mixture distribution defined for $\mathcal{D}$. Saeed et al. (2020) provides an auxiliary graphical representation, the mixture DAG (Definition 3.1), to easily identify d-separation for variable with respect to a collection DAGs.
The mixture DAG, denoted $\mathbf{D}_{\mathcal{M}}$, has vertex set $\mathbf{V}_{\mathcal{M}}=\left(\bigcup_{l=1}^{k}\mathbf{V}_{l}\right)\cup\{d\}$, defined as the union of all the vertices of the $k$ component DAGs in the set $\mathcal{D}=\{\mathbf{D}_{1},\dots,\mathbf{D}_{k}\}$ and an additional vertex, $d$, for the DAG. Let, $i_{l}$, be the $i$th vertex of DAG $l$ and $[i]$ denote the set $\{i_{1},\dots,i_{k}\}$. The edge set, $\vec{\mathbf{E}}_{\mathcal{M}}$ consists of the directed edges $\vec{\mathbf{E}}_{l}$ for $l=1,\dots,k$ and additional directed edges from the vertex $d$ to each vertex $i_{l}$ for which the parents of $i$ differ in at least one of the DAGs in $\mathcal{D}$. By definition, if $(i_{h},d)\in\vec{\mathbf{E}}_{\mathcal{M}}$ for one of the component DAGs, then we have $(i_{l},d)\in\vec{\mathbf{E}}_{\mathcal{M}}$ for $l=1\dots,k$.
Now, with this auxiliary mixture DAG, we can make statements about the conditional independence sets of vertices and their associated variables. Let $[A]$ be the set of vertices, $\{A_{1},\dots,A_{k}\}$, across the $k$ component DAGs. If $[A]\perp[B]\left|[C]\right.$ in $\mathbf{D}_{\mathcal{M}}$ by the rules of d-separation, then $\mathbf{z}_{A}\perp\mathbf{z}_{B}|\mathbf{z}_{C}$ in the mixture distribution defined on $\mathcal{D}$ (Theorem 3.2, Saeed et al., 2020).
In a standard MDGM model, the parents of each vertex will vary across the collection of DAGs used in the mixture. This is a desirable probability so that the directional dependence is not encoded a priori into the model. As a consequence, $\mathbf{z}_{A}\not\perp\mathbf{z}_{B}|\mathbf{z}_{C}$ for any $A$, $B$ and $C$ in the MDGM. This indicates that an MDGM, even with strongly compatible DAGs, cannot strictly represent the same conditional independence relationships as an MRF. The dependence, however, arises solely from the variables’ dependence on the DAG structure. The conditional independence relationships described for the DAGs above still hold when the DAG is included in the conditioning set. Conditioned on the DAG, strongly compatible DAGs will not introduce dependencies which are not already present in the NUG.
Appendix H Simulation Study
H.1 Phase Transition
@@
H.2 Complete Data Setting
In the complete data scenario, we set $m_{i}=2$ for all $i=1,\dots,n$ and perform the simulation across all combinations of $\beta=\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8\}$ and $\eta=\{0.05,0.2\}$. The results are shown in Figure 8.
@@ H.2 Complete Data Setting
Consistent with the missing data regime, we observe similar performance of all five models in terms of accuracy and RMSE for the low uncertainty setting (i.e. $\eta=0.05$). Likewise, in the high uncertainty setting, the aMRF has noticeably worse performance in estimating the true clustering of the latent field as the spatial dependence parameter increases, while the MDGM-AO and MDGM-ST come close to matching the performance of the MRF. The aMRF shows slightly worse performance in terms of accuracy as $\beta$ increases in the high uncertainty setting, with the MDGM-AO class performing worse as well when $\beta=0.8$. In terms of posterior approximation, again the MDGM-ST demonstrates the best performance, particularly in the high uncertainty setting.
H.3 $32\times 32$ Lattice
Figure 9: Simulation study results for the complete data setting on a $32\times 32$ grid. The points are the estimated expected value of the statistic with error bars giving the 90% bootstrap confidence interval of the Monte Carlo error.
Figure 10: Simulation study results for the missing data setting for a $32\times 32$ grid. The points are the estimated expected value of the statistic with error bars giving the 90% bootstrap confidence interval of the Monte Carlo error.
For the larger lattice (four times as many areal units), we compare the MDGM-ST to the aMRF for the high uncertainty complete and missing data regimes. We followed the same initialization procedure as described in Section 5, obtaining five thousand posterior draws and discarding the first one thousand as burn-in. Only two fitted posteriors across all simulations were thrown out due to indications of lack of convergence (R-hat $>1.1$). Figures 9 and 10 show the bootstrap confidence intervals for the expected posterior accuracy and posterior RMSE. Both models perform well in terms of accuracy. Performance decreases relative to the MDGM-ST in estimating the true spatial dependence for the aMRF in both the complete and missing data settings, but with closer performance in both settings when $\beta=0.8$. On the 4th generation Intel processor, the MDGM-ST averaged 2.4 hours for five thousand draws, while the aMRF took 14.5 minutes on average. On the 2nd gen. processor, the MDGM-ST took 3.6 hours and the aMRF 27.7 minutes on average. Our implementation of Wilson’s algorithm did not quite reach a linearithmic ratio between the run times, but is still subquadratic. Run time performance can be substantially improved by implementation in a lower level language such as C++.
Appendix I Cross Validation Study
For each replication of the cross validation study we randomly select 60 areal units with ratings to be withheld as the test set. Using the remaining ratings as our training set, we fit the model with both a MDGM-ST and aMRF prior on the latent variable. As the withheld data is the observed ratings, we compare using the mean absolute difference of the posterior mean predicted probability of a block group having a rating of one to the average observed ratings at the block group. We take the average of the absolute errors for the block groups in the test set as our summary statistic of model performance for a single replication of the study. We performed one hundred replications of the cross validation study. We found that the mean absolute error of the posterior predicted probability of a rating of one to the average observed ratings for the model with an MDGM-ST prior is .2188, and the mean absolute error for the aMRF prior is .2204. The maximum absolute difference between the mean absolute errors of the two models across replications was .007, indicating a practically equivalent performance of the two models.